OpenStudy (anonymous):

Solve for t. 2ln|t + 4| = 1 Please & thank you :) Absolute value t + 4 btw if you couldn't tell!

OpenStudy (anonymous):

first use this property \[aln(b)=ln(b^a)\]

OpenStudy (anonymous):

so just ln|t+4|^2.

OpenStudy (anonymous):

alright now use \[e^{ln(b^a)}=b^a\]

OpenStudy (anonymous):

but remember if you make e the base on the left side you must do it on the right side

OpenStudy (anonymous):

What would the answer be...this is one of the problems on my AP calc summer assignment & I don't remember this from precalc

OpenStudy (anonymous):

well you'd get |t+4|^2=e^1

OpenStudy (anonymous):

-2.35?

OpenStudy (anonymous):

jk....

OpenStudy (anonymous):

not quite take the square rt of both

OpenStudy (anonymous):

I did, so it would be |t+4| = √e

OpenStudy (anonymous):

\[|t+4|=\sqrt{e}\]

OpenStudy (anonymous):

Then subtract 4

OpenStudy (anonymous):

yes however you have the variable t so you should be + -

OpenStudy (anonymous):

Okay, that's because of the absolute value right?

OpenStudy (anonymous):

wait it will always be positive neverm

OpenStudy (anonymous):

Okay lol thank you

OpenStudy (anonymous):

so basically t+4=-sqrt{e} and t+4=sqrt{e}

OpenStudy (anonymous):

Thanks again!

OpenStudy (anonymous):

\[\ln|t+4|=\frac{1}{2}\] \[|t+4|=e^{\frac{1}{2}}\] \[t+4=e^{\frac{1}{2}}\] or \[t+4=-e^{\frac{1}{2}}\]